Boatler Used Cadillac Co. requires $930,000 in financing over the next two years. The firm can borrow the funds for two years at 10 percent interest per year. Ms. Boatler decides to do forecasting and predicts that if she utilizes short-term financing instead, she will pay 6.75 percent interest in the first year and 11.55 percent interest in the second year. Assume interest is paid in full at the end of each year a. Determine the total two-year interest cost under each plan. Interest Cost Long-term fixed-rate Short-term variable-rate b. Which plan is less costly? Short-term variable-rate plan Long-term fixed-rate plan
Answer #1

a.

Interest cost in long term fixed- rate = 930,000 * 10% * 2 = 186,000

Interest cost in short term variable = Int. for first year + Int. for second year = 930,000 * 6.75% + 930,000 * 11.55% = 170,190

b.

Short-term variable rate is cheaper.